# Representative Action Selection for Large Action Space Meta-Bandits

### v2

@@ 2.2 Epsilon Net Algorithm
We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by Haussler and Welzl (1986). It selects $K$ i.i.d. random actions, aligned with the distribution of bandit instances. Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
@@
7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$triangleright Repetition of actions is allowed
@@
The optimal action $a^{*}(\theta)$ is unique with probability 111 for all bandit instances.
@@
[…]
@@
By definition, Algorithm 1 samples $K$ i.i.d. extreme points from clusters in $\mathcal{R}$ according to measure $q$. If a cluster $r\in\mathcal{R}$ has a higher measure $q(r)$, its elements are more likely to be included in the output. In fact, with high probability, this algorithm outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$ with respect to measure $q$. (This is a simplified version of Theorem 10.2.4 of Matousek (2013).)
@@
The partition $\mathcal{R}$ bridges the two definitions of $\epsilon$-net: choosing one point from each cluster gives an $\epsilon$-net in both senses, though with different values of $\epsilon$. Geometrically, $\epsilon$ is the largest cluster diameter $\epsilon:=\max_{r\in\mathcal{R}}\mathrm{diam}(r)$; measure-theoretically, $\epsilon$ is the smallest cluster measure $\epsilon:=\min_{r\in\mathcal{R}}q(r)$. Hence, we have the following corollary:
Corollary 5.
Under the same conditions as Lemma 4, let $\epsilon:=\min_{r\in\mathcal{R}}q(r)$. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, the following holds:
$\forall a\in\mathcal{A}_{\operatorname*{full}},\ \exists a^{\prime}\in\mathcal{A}\text{ such that }\|a-a^{\prime}\|_{2}<\max_{r\in\mathcal{R}}\operatorname{diam}(r).$
@@ 3 Regret Analysis
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$, and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
We impose the following assumption on the partition, used only for lower bounds: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it. This ensures clusters are well-separated.
Assumption 6 (Effective partition).
@@
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, there is an absolute constant $C>0$, such that
@@ 3.1 Regret Bounds of Algorithm
The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds. The key difference is that whereas the reference subset includes a representative from each cluster, the algorithm may miss some clusters. However, although it may select a smaller subset, the algorithm achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
@@
Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be output of Algorithm 1. For the same constant $C>0$ in Theorem 7,
Proof sketch. If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, choose $a_{\ell}\in r_{\ell}\cap\mathcal{A}$ as the representative point. If $r_{\ell}\cap\mathcal{A}=\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$. The new set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset. Then, for each $\ell\leq m$, define a simple Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, we consider two cases:
  • $r_{\ell}\cap\mathcal{A}\neq\emptyset$: The regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$.
  • $r_{\ell}\cap\mathcal{A}=\emptyset$: The regret is bounded by $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\!\mu_{a}$. $\square$
The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 8 behaves similarly to entropy: it is maximized when $q$ is uniform, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Lemma 18 in Supplementary. Intuitively, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$. This expression is small when $q(r)$ is small, and decays faster when $q(r)$ gets larger, making the overall term negligible if $q$ is highly concentrated.
The connection between Theorem 7 and 8 is insightful:
@@
On the other hand, if the measure $q$ is close to uniform, then $\min_{r\in\mathcal{R}}q(r)$ is large. In this case, with relatively high probability, Algorithm 1 outputs a geometric $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$-net (see Corollary 5). Moreover, since the partition is constructed using a grid, the cluster diameters are similar, so $\max_{r\in\mathcal{R}}\mathrm{diam}(r)$ closely matches the spacing of the original grid—making the output resemble a slightly coarser grid.
Partition-Independent Upper Bound.
@@
$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big{\{}m\!\in\!\mathbb{N}:$$\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},$
$\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big{\}}.$
@@
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;$ $2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},$
$\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
@@
[…]
Lower Bound.

### v3

@@ 2.2 Epsilon Net Algorithm
We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by haussler1986epsilon . It selects $K$ i.i.d. random actions, aligned with the distribution of bandit instances. Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
@@
7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$ $\triangleright$ Repetition of actions is allowed
@@
The optimal action $a^{*}(\theta)$ is unique with probability $1$ over all bandit instances.
@@
where $p(\theta)$ is the density of $\theta$ and $\int p(\theta)d\theta=1$. Under Assumption 2, measure $q$ is a probability distribution. It reflects the probability that a given cluster contains the optimal action and thus represents the potential contribution of that cluster to the expected regret.
@@
By definition, Algorithm 1 samples $K$ i.i.d. extreme points from clusters in $\mathcal{R}$ according to measure $q$. If a cluster $r\in\mathcal{R}$ has a higher measure $q(r)$, its elements are more likely to be included in the output. In fact, with high probability, this algorithm outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$ with respect to measure $q$. (This is a simplified version of Theorem 10.2.4 of matousek2013lectures .)
@@
The partition $\mathcal{R}$ bridges the two definitions of $\epsilon$-net: choosing one point from each cluster gives an $\epsilon$-net in both senses, though with different values of $\epsilon$. Geometrically, $\epsilon$ is the largest cluster diameter $\epsilon:=\max_{r\in\mathcal{R}}\mathrm{diam}(r)$; measure-theoretically, $\epsilon$ is the smallest cluster measure $\epsilon:=\min_{r\in\mathcal{R}}q(r)$.
@@ 3 Regret Analysis
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$ and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
Only for lower bounds, we assume a well-separated partition: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it.
Assumption 5.
@@
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, there is an absolute constant $C>0$, such that
@@
For each $\ell\leq m$, the term $\mathbb{E}_{\theta}[\max_{a\in r_{\ell}}\mu_{a}]$ equals $\mathbb{E}_{\theta}[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}]$, since $\mathbb{E}_{\theta}[\mu_{a}]=0$.
@@ 3.1 Regret Bounds of Algorithm
The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds. The key difference is that whereas the reference subset includes a representative from each cluster, the algorithm may miss some clusters. However, the algorithm still achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
@@
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be output of Algorithm 1. For the same constant $C>0$ in Theorem 6,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof sketch. If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, choose $a_{\ell}\in r_{\ell}\cap\mathcal{A}$ as the representative point. If $r_{\ell}\cap\mathcal{A}=\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$. The new set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset. Then, for each $\ell\leq m$, define a Gaussian process $\{Z_{a}\}_{a\in r_{\ell}}$, where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$, and let $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, we consider two cases. If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, the regret is upper bounded by $Y_{\ell}$, and hence by $\max_{\ell\leq m}Y_{\ell}$. If $r_{\ell}\cap\mathcal{A}=\emptyset$, the regret is bounded by $\max_{a\in\mathcal{A}_{\text{full}}}\mu_{a}$. $\square$
The term $\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]$ in Theorem 7 is maximized by the uniform $q$, provided $|\mathcal{R}|\geq 2K+1$ (since we can choose any partition); see Supplementary. Intuitively, each cluster $r$ contributes $q(r)\cdot(1-q(r))^{2K}$. This expression is small when $q(r)$ is small, and decays faster when $q(r)$ gets larger, making the overall term negligible if $q$ is highly concentrated.
The connection between Theorem 6 and 7 is insightful:
@@
$N(\mathcal{A}_{\operatorname*{full}},\!\epsilon)=\min\Big\{m\!\in\!\mathbb{N}:\;\exists\{a_{\ell}\}_{\ell\leq m}\subseteq\mathcal{A}_{\operatorname*{full}},\forall a\in\mathcal{A}_{\operatorname*{full}},\exists\ell,\|a\!-\!a_{\ell}\|_{2}\leq\epsilon\Big\}.$
@@
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\leq\;2\epsilon\sqrt{n}+C\epsilon\sqrt{\log N(\mathcal{A}_{\operatorname*{full}},\epsilon)},\text{where }K\geq$ $c\cdot(M/\epsilon)^{2}\cdot N(\mathcal{A}_{\operatorname*{full}},\epsilon).$
@@
The partition-independent upper bound in Theorem 9 depends on a positive constant $\epsilon$, which serves as a tolerance parameter up to the diameter of the action space. Given a $\epsilon$, the required number of samples $K$ scales with the square of the diameter-to-$\epsilon$ ratio, multiplied by the covering number. The resulting expected regret is then bounded in terms of $\epsilon$, the dimensionality, and the logarithm of the covering number. As $\epsilon$ decreases, more samples are needed, but the regret bound becomes tighter.
The matching lower bound for Theorem 7, as well as the upper bound under the general sub-Gaussian process assumption for Algorithm 1, are provided in the Supplementary.
4 Generalization and Empirical Validation
In general cases, the decision-maker may not have explicit access to the full structure of the action space, especially in high-dimensional settings. Instead, they are given a list of actions and can observe the expected outcomes of these actions through sampling. We therefore treat the expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ as a family of random variables indexed by an abstract set, e.g., $\mathcal{A}_{\operatorname*{full}}:=\{\textnormal{treatment A},\textnormal{treatment B},\textnormal{treatment C}\}$. In a single bandit, these expected outcomes define an outcome function $f$ over the action space $\mathcal{A}_{\operatorname*{full}}$: $f(a):=\mu_{a},\forall a\in\mathcal{A}_{\operatorname*{full}}$. We show that our framework applies when the outcome functions lie in a reproducing kernel Hilbert space (RKHS) (williams2006gaussian, ). Consider a positive semidefinite kernel $k:\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}\to\mathbb{R}$. By Mercer’s theorem, for a non-negative measure $\mathbb{P}$ over $\mathcal{A}_{\operatorname*{full}}$, if the kernel satisfies $\int_{\mathcal{A}_{\operatorname*{full}}\times\mathcal{A}_{\operatorname*{full}}}k^{2}(a,a^{\prime})d\mathbb{P}(a)d\mathbb{P}(a^{\prime})<\infty,$ then it admits an eigenfunction expansion: $k(a,a^{\prime})=\sum\nolimits_{i\leq\infty}\lambda_{i}\phi_{i}(a)\phi_{i}(a^{\prime}),$ where $(\phi_{i})_{i\leq\infty}$ are orthonormal eigenfunctions under $\mathbb{P}$, and $(\lambda_{i})_{i\leq\infty}$ are non-negative eigenvalues. Let outcome functionff be of the form $f(\cdot)=\sum_{i=1}^{N}\alpha_{i}k(\cdot,a_{i})$ for some integer $N\geq 1$, and a set of points $\{a_{i}\}_{i=1}^{N}\subset\mathcal{A}_{\operatorname*{full}}$ and a weight vector $\alpha\in\mathbb{R}^{N}$. The function $f$ can be rewritten as
$\mu_{a}=f(a)=\langle\mathbf{f},\Phi(a)\rangle,$
where $\mathbf{f}$ is a vector of coefficients, and $\Phi(a)$ usually refereed as a feature map, has entries $\Phi_{i}(a)=\sqrt{\lambda_{i}}\phi_{i}(a)$. The full action space is the one formed by the feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{\operatorname*{full}}$.
Assuming the outcome function is generated from a kernel inherently implies that actions are correlated. To study the effect of varying correlation levels among actions, we use the parametric RBF kernel to sample outcome functions from a RKHS: $k(a,a^{\prime}):=\exp\left(-\frac{\|a-a^{\prime}\|^{2}}{2l^{2}}\right)$ , where $l$ is a length-scale parameter that control the dependence among actions. By (kanagawa2018gaussian, , Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$. Note that although the kernel is used to generate outcome functions, the kernel itself is not revealed; consequently, $a$ and $a^{\prime}$ are regarded merely as indices rather than as inputs to the kernel.
Figure 1: Comparison on solving equation 5, selecting $K=5$ distinct actions from 15 grid points over $[0,2]$, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales. Left: Expected regret over 50 repetitions of our method with exhaustive search (green), compared with Thompson Sampling (TS; yellow), Upper Confidence Interval (UCB; orange), and Successive Halving (SH, blue) applied to the set of super-arms. TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls. Right: SH’s expected regret (blue) and pull counts (yellow bars) per round, compared with our method (green). SH requires nearly 10,000 pulls—about three times the number of super-arms—to match our method.
Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in equation 1. Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
In general, the problem is non-trivial since the objective involves an expectation over random variables with an unspecified distribution, and the inner expression is a maximum over a subset of potentially correlated variables. Consequently, small changes in $\mathcal{A}$ can cause non-smooth variations in the maximum. Our method is computationally efficient, avoiding the combinatorial complexity.
We assess performance in identifying near-optimal solutions by selecting $K=5$ actions from an action space of 15 grid points in $[0,2]$. The configuration yields a moderate number of super arms, allowing approximation of the optimal subset using Successive Halving (SH), which, however, becomes impractical for larger action spaces. Outcome functions are generated with an RBF kernel at length-scales $l\in\{0.5,1,\dots,4\}$. We compare the proposed method against Thompson Sampling (TS) and Upper Confidence Bound (UCB). In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where exhaustive search is used over the action space to find the optimal action $a^{*}(\theta)$. For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$. Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective. In TS/UCB methods, we adopt a bandit feedback setting: at each round (corresponding to a bandit instance $\theta$), the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}(\theta)$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$). Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB, with the prior set to $\mathcal{N}(0,1)$. In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in karnin2013almost , $N\log_{2}N\approx 37{,}000$) over a few rounds. In each round (corresponding to a bandit instance $\theta$), the number of remaining super arms is halved. This process continues until only the best-performing super arm remains or the budget is exhausted. We evaluate expected regret of all methods over $10^{5}$ additional randomly sampled outcome functions.
The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions. As shown, the expected regret decreases as the length-scale increases. The length-scale $l$ controls the effective number of approximately independent actions: when $a$ and $a^{\prime}$ are far relative to $l$, $k(a,a^{\prime})$ is negligible, so $\mu_{a}$ and $\mu_{a^{\prime}}$ are nearly uncorrelated. In the limit, $\{\mu_{a}\}_{a\in\mathcal{A}_{\text{full}}}$ forms a collection of i.i.d. zero-mean, unit-variance variables.
The SH in left subplot of Figure 1 is used to approximate the ground-truth optimal subset and our method does give similar performance as SH. The right subplot highlights SH’s impracticality: we ran our method (green) and SH (blue) for 50 repetitions with fixed $l=1$, tracking the current best super arm and total super arm pulls after each SH round. The blue curves and shades represent the mean $\pm$ one standard deviation of SH’ expected regret, while yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round. Green curves show our method’s expected regret. SH matches our performance after three rounds but requires 9005 pulls—roughly three times the number of super arms—making it intractable for large $\mathcal{A}_{\operatorname*{full}}$.
Figure 2: Comparison of our method (green), CTS (brown), and CUCB (red) on solving equation 5 with $K=10$ actions chosen from 500 grid points over [5,5][-5,5], with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales. In our method, EpsilonNet+TS, the “argmax” step in Algorithm 1 is replaced by TS run for 300 rounds. CTS is run for 3000 rounds and CUCB for 2000 rounds. All methods use prior variance 11. Left: expected regret over 30 repetitions. Right: runtime (purple, left axis) and number of arm pulls (gray, right axis) over 30 repetitions.
Further, we demonstrate the generality of our method (denoted EpsilonNet+TS), where the “argmax” step in Algorithm 1 is replaced by TS. We consider the same setup as in Figure 1, but with $K=10$ actions chosen from 500 grid points over [5,5][-5,5]. In our method, for each bandit instance $\theta$, we run TS for 300 rounds, selecting an action $a$ and observing $\mu_{a}$ in each round; the final action is taken as an approximation of the optimal action $a^{*}(\theta)$. We repeat this process until $K=10$ distinct actions are identified. We compare against two heuristics: combinatorial TS (CTS) and combinatorial UCB (CUCB). CTS is run for 3000 rounds to match the runtime of our method, while CUCB is run for 2000 rounds due to its slower execution and additional initialization step. Both CTS and CUCB operate in the semi-bandit feedback setting: in each round (corresponding to one bandit instance $\theta$), an outcome function is sampled and $K=10$ actions $\mathcal{A}$ are selected, with the rewards $\mu_{a}(\theta)$ revealed for all $a\in\mathcal{A}$. The prior for TS, CTS, and CUCB is set to $\mathcal{N}(0,1)$, consistent with the ground truth.
In Figure 2, the left subplot reports the expected regret and its standard deviation for EpsilonNet+TS (green), CTS (brown), and CUCB (red), averaged over 30 repetitions. The poor performance of CTS and CUCB is expected, as noted in related works. The variance of EpsilonNet+TS could be further reduced by running more TS rounds or using a more accurate action-identification method, since we currently use a Gaussian prior—modeling each action’s expected outcome $\mu_{a}$ as an independent Gaussian random variable—rather than a Gaussian process prior, in order to accommodate the most general case where $\mathcal{A}_{\operatorname*{full}}$ is an index set. In the right subplot, we fix the length-scale $l=1$ and evaluate runtime and arm pulls of the three methods over 30 repetitions. The left axis (purple) shows average runtime $\pm$ one standard deviation, and the right axis (gray) shows the corresponding number of arm pulls. EpsilonNet+TS requires 3000 pulls, since it selects $K=10$ optimal actions, with each action identified by running TS for 300 rounds. CTS has similar runtime but needs far more pulls, as it runs 3000 rounds with $K=10$ actions pulled per round.
Conclusion and Future Work
We proposed a framework for selecting a subset of correlated actions. A simple algorithm was introduced and shown to effectively identify near-optimal subsets. Future work could develop a stopping criterion. When subset size is flexible but sampling new bandit instances is costly, ideas from species discovery (roswell2021conceptual, ) may be useful—treating the discovery of a new action as analogous to discovering a new species.
Appendix Summary: The appendix is divided into four parts.
First, Sections A-B provide additional related works and the regret bound of Algorithm 1, which were omitted from the main body due to page limits.
Second, Sections C to H contain the proofs of all the theorems and lemmas stated both in the main body and in Section B.
Third, Sections I-J present auxiliary tools used throughout the paper.
Fourth, Sections K-M include additional examples and a numerical study of Algorithm 1. This part features a low-dimensional ($n=3$) action space on the unit sphere to investigate the effect of cluster diameter on expected regret, in line with Theorem 6. It also studies two extensions of the varying dependence structure of actions introduced in the main body:
  • Nonstationary dependence over the action space: Using a nonstationary Gibbs kernel, where the kernel value depends not only on the difference between two inputs $a,a^{\prime}$ but also on the location of $a$.
  • Limiting case: The collection of expected outcomes $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ becomes i.i.d.
LLM Usage: Large language models are used for polishing writing, and finding relevant research.
Appendix A Extra Related Works
Epsilon Nets have two standard definitions. The first, geometric definition [37], requires that radius-$\epsilon$ balls centered at net points cover the set. It relates to the covering number and extends to function classes, as in [31]. The second, measure-theoretic definition [24], requires the net to intersect all subsets of sufficiently large measure. The classic $\epsilon$-net algorithm by [14] remains the simplest and most broadly applicable method. Later works aim to reduce net size [26, 27, 25] and address online settings [6].
Expected supremum of Gaussian process for a given set $\mathcal{S}$ refers to the term $\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]$. It is an important topic in high-dimensional probability [37]. The sharpest known bounds are due to [34]. and asymptotic convex geometry [29, Chapter 9]
Appendix B Extra Regret Bounds of Algorithm 1
We give a matching lower bound: